# RPTS: Tree-Structured Reasoning Process Scoring for Faithful Multimodal Evaluation

### v1

@@ Abstract
Codehttps://aaai.org/example/code
Datasetshttps://aaai.org/example/datasets
Extended versionhttps://aaai.org/example/extended-version
@@ Introduction
Recent advances in multimodal foundation models have demonstrated increasingly sophisticated capabilities in the combination of visual and textual information  (gpt4). However, as these models begin to assist in evidentiary reasoning tasks such as criminal case analysis - where establishing reliable connections between surveillance footage (visual modality), forensic reports (textual modality), and other evidence is crucial, and where conclusions must follow rigorous, verifiable reasoning chains - two critical questions emerge: 1. Can current evaluations distinguish between logically valid reasoning and coincidentally correct conclusions? 2. Do existing frameworks capture the non-linear, cross-modal reasoning required to resolve conflicting evidence?
@@
Most existing benchmarks focus solely on task accuracy through multiple choice or short answer formats  (mmvet; mmmu), completely ignoring the reasoning process. This approach fails to detect when models arrive at correct conclusions through flawed reasoning, a phenomenon we call ”right answers for wrong reasons” , as illustrated on the left side of Figure 1. The few works that examine reasoning processes  (roscoe; receval) typically adopt an oversimplified linear evaluation framework. These approaches are fundamentally mismatched to the complex, non-linear nature of real-world reasoning, where multimodal evidence may appear conflicting yet collectively support valid conclusions. To address these limitations, we introduce a novel evaluation metric: Reasoning Process Tree Score (RPTS), designed to assess multimodal reasoning processes. The core innovation of RPTS lies in its tree-structured representation of reasoning, where leaf nodes correspond to atomic evidence units (visual or textual) and non-leaf nodes capture the hierarchical derivation of intermediate conclusions. This structure inherently accommodates the non-linear interactions characteristic of multimodal reasoning. Furthermore, RPTS incorporates two key hyperparameters whose adjustable values enable precise quantification of both global and local logical consistency, thereby facilitating accurate error localization within the reasoning chain.
@@
Classic multimodal benchmarks typically assess the specific reasoning abilities of the models. For example, OK-VQA(okvqa) evaluates a model’s capacity to leverage external knowledge for reasoning, while VCR(vcr) focuses on human-related common sense reasoning. To evaluate the comprehensive capabilities of a model, researchers have proposed various benchmarks, such as MMBench(mmbench), SEED-Bench(seedbench), MM-Vet(mmvet), and MMMU(mmmu). These benchmarks scrutinize the reasoning abilities of models from diverse perspectives, often employing multiple choice or simplified formats to facilitate the evaluation process. InfiMM-Eval(infimm-eval) incorporates the reasoning process into the evaluation, scoring the entire reasoning process. However, it cannot perform a more detailed analysis of reasoning and its evaluation method cannot exclude cases where incorrect reasoning leads to a correct answer.
@@
Recent studies have introduced various techniques for evaluating reasoning processes. ROSCOE(roscoe) proposes a set of quality metrics to assess reasoning from four perspectives: semantic alignment, semantic similarity, logical correctness, and semantic coherence. ReCEval (receval) evaluates reasoning based on two criteria: whether the reasoning steps are correct and whether new information is derived from the reasoning. REVEAL provides a dataset to validate whether a model can be used to verify the reasoning process.
@@
To construct the reasoning tree, we first parse the model’s reasoning into a structured format: ”[PREMISE] + [PREMISE] $\rightarrow$ [CONCLUSION]”, where ’[PREMISE]’ can be derived from visual clues, textual clues, or intermediate conclusions from prior steps. However, existing open-source MLLMs cannot strictly adhere to this output format. To address this, we first employ chain-of-thought (CoT) prompting to guide the model in generating step-by-step reasoning with explicit premises. Subsequently, we use GPT-4 to reformat the reasoning into a structured, easily parsable representation. The parsed reasoning results can also be utilized for accuracy computation.
@@
Now, each reasoning step in our approach strictly adheres to the ”[PREMISE] + [PREMISE] $\rightarrow$ [CONCLUSION]” format. Prior studies (llm-evaluator; G-Eval; gptscore; touchstone; visit-bench; mmvet; infimm-eval) have demonstrated LLMs’ effectiveness in assessing model reasoning. Therefore, we utilize a LLM to score reasoning, but with a unique twist: we only evaluate individual reasoning steps, not the entire process. This method allows for more precise evaluations by preventing the influence of other reasoning elements on the scores. Before we input the reasoning into scorer, we first preprocess the model’s reasoning by eliminating redundant text clues, merging conclusions from images, substituting unnumbered texts and conclusions with all relevant clues and conclusions, and removing reasoning without ’[PERMISE]’. For scoring reasoning according with image, we calculate the semantic similarity of conclusions directly derived from images against the ground truth. For other reasoning, we input the premises and conclusion into LLM to assess their logical coherence. The score given by scorer ranges from 0 to 1, with higher scores indicating stronger logical reasoning. However, as illustrated in Figure 1, there are instances where the model’s selected premises may not directly support the given conclusion, though they may be justified within the broader reasoning context. To address this, if the initial score is below 0.5, we re-evaluate using all text clues and previously derived conclusions as new premises, and then applying a 0.8 penalty for incorrect premises. We select the higher of the two scores as the final assessment.
@@
Figure 5: Three types of relationships between modalities. The red arrow means the relation between image and text, and the blue arrow means the interference between them.
@@
Figure 4 illustrates the scoring outcomes under three different settings of $\lambda$ and $h_{f}$. If we aim to assign equal importance to each reasoning step, we can set $\lambda$ to 1, as shown in the top-right corner of Figure 4. In this case, RPTS represents the average score of all reasoning steps. Conversely, if we only wish to focus on the first step, i.e., reasoning at height 1, as depicted in the bottom-left corner, we set $\lambda$ to 0 and $h_{f}$ to 1. Under this configuration, only the score of the first step contributes to the RPTS computation. In the main experiment presented in Table 4, the hyperparameters we configured are consistent with those shown in the bottom-right corner of Figure 4. We focused more on the first step in reasoning, so we set $h_{f}=1$. As shown in Table 1, most of the reasoning trees in RPTS-Eval have a height of no more than 7, and we want the weight of each reasoning step to be no less than 0.5 when calculating the RPTS. Based on Formula 1, we can calculate $\lambda>\sqrt[|h_{f}|-7]{0.5}\approx 0.891$. Therefore, we selected $\lambda=0.9$.
@@
To reduce the difficulty of manually designing stories, we use GPT-4 to assist annotators. First, we ask an annotator to design a few reasoning stories and input them into GPT-4 as examples. Following the approach of MM-Vet(mmvet), we then require GPT-4 to generate reasoning stories encompassing six types of capabilities, based on the given examples. However, we define two distinct capabilities that differ from MM-Vet: Image Comparison (IC) and Spatial Awareness (SA). The remaining four capabilities—Recognition (Rec), OCR, Commonsense Reasoning (Com), and Math—are identical to those defined in MM-Vet. The specific definitions of IC and SA are as follows:
@@
As mentioned above, the stories generated by GPT-4 lack logical consistency. Therefore, annotators only draw inspiration from these stories rather than using them directly, thereby reducing the difficulty of story design. For example, GPT-4 generates a story set on a rainy day, and our annotators draw inspiration from this, such as the idea of ’rain,’ and design reasoning tasks based on that inspiration. For instance, they might infer whether the box is open or closed based on whether there’s water inside, or they might reason about whether outdoor activities can continue based on certain pre-rain features. The six capabilities are as follow:
@@
To ensure data quality, each piece of data is validated by two validators. We reference InifMM-Eval(infimm-eval) and conduct a comprehensive evaluation of the data based on the following criteria:
@@
Benchmark Size Images Answer Format Metric Evaluate Reasoning
MMMU(mmmu) 11.5K 12.5K Option/Open Answer Accuracy
$M$-Vet(mmvet) 218 200 Shot answer GPT4-score
InifMM-Eval(infimm-eval) 279 342 Reasoning GPT4-score
RPTS-Eval(Ours) 390 374 Reasoning Accuracy+RPTS
Experiments
Models and Evaluation Metrics
To validate the challenging nature of RPTS-Eval and the capability of the RPTS evaluation metric analysis model, we conducted experiments in both Chinese and English across various models. The open-source models tested include InstructBLIP(instructblip), InternVL2(intervl2), ShareGPT4V(sharegpt4v), Llava-v1.5(llava1.5), Llava-Next(llavanext) and Qwen-VL-Chat(qwenvl), detailed in Appendix A; the sole close-source model examined is GPT-4o. We evaluate the reasoning ability of the model by combining accuracy and RPTS, and analyze the problems of the model.
Scorer Selection
Model $\overline{\Delta}$
Qwen2-7B 0.216
Llama-3-8B 0.231
Qwen2-72B 0.143
Llama-3-70B 0.152
GPT-4 0.095
Experiment Settings
Our experiment involves both Chinese and English languages and performs chain-of-thought(COT)(cot) reasoning on the RPTS-Eval benchmark. All tests were performed in a zero-shot setting using a greedy decoding strategy to assess the models’ inferential abilities. To optimize the COT reasoning outcomes, we designed five Chinese prompts and seven English prompts, selecting the most effective one from each language for our experiments. All tests were carried out on an NVIDIA A100 GPU. When reasoning, we set the temperature of each model to 0 and use greedy decoding.
@@
Table 4 displays the performance of various models on RPTS-Eval. In addition to evaluating the accuracy of the models’ inferences and their mean RPTS scores, we applied an RPTS-based filter to exclude cases where incorrect inferences resulted in accurate conclusions. Specifically, we consider the reasoning logic with an RPTS score below 0.5 to be incoherent, and therefore classify it as incorrect. The data in Table 4 reveal that all models exhibited a decline in accuracy to varying extents, with GPT-4o showing the least reduction. This modest decline is closely associated with GPT-4o’s advanced logical capabilities. In the results of GPT-4, the lower RPTS scores are associated with erroneous reasoning and the model’s failure to capture certain infomation. Conversely, the open-source models demonstrated a lack of logical robustness in their reasoning processes, leading to more pronounced decreases due to often generating irrelevant or illogical outputs. Despite these models’ lower accuracy, their RPTS scores were not significantly impacted. We hypothesize that this is due to two primary reasons: 1. Disconnection between the inference outcomes and the intended targets. While the models initially could reason based on the specified targets, they gradually lost focus on the targets as the number of reasoning steps increased, resulting in conclusions that diverged from the intended data targets. 2. Recurrent generation of identical sentences. Across various sizes, the open-source models consistently produced repetitive reasoning that, while logically sound, failed to reach the desired conclusions. These factors led to reduced accuracy but did not substantially affect the logical integrity of the inferences, as reflected in the relatively high RPTS scores. In addition, Appendix B shows the performance of the model on six capabilities.
Step Analysis.
@@
Image Text Image Text
IV-8B 0.50 0.76 0.54 0.94
Lv-7B 0.42 0.78 0.40 0.67
LN-7B 0.36 0.52 0.22 0.58
SG-7B 0.35 0.62 0.30 0.70
Lm-11B 0.52 0.83 0.4 1.0
IB 0.40 0.69 - -
Lv-13B 0.41 0.66 0.37 0.78
LN-13B 0.36 0.57 0.29 0.73
Qwen 0.45 0.74 0.45 0.66
SG-13B 0.18 0.57 0.42 0.75
IV-26B 0.53 0.80 0.57 0.84
LN-34B 0.54 0.80 0.52 0.72
IV-40B 0.61 0.90 0.60 0.88
IV-76B 0.60 0.92 0.60 0.87
Lm-90B 0.58 0.79 0.6 0.75
@@
Table 5: RPTS score for drawing conclusions from visual clues or textual clues. IB: InstructBLIP; IV: InternVL2; Lm: Llama-3.2; Lv: Llava-v1.5; LN: Llava-Next; Qwen: Qwen-VL-Chat; SG: ShareGPT4V;
Sensitivity Analysis
@@
Table 6: A sensitivity analysis for RPTS metric. The upper table displays the RPTS values under different configurations of $\lambda$ and $h_{f}$. The lower table presents the percentage of reasoning paths, among the total, where the answer is correct but the RPTS value is less than 0.5, across various $\lambda$ and $h_{f}$ settings.

### v2

@@ Abstract
Code & Datasetshttps://github.com/wang-hao-feng/RPTS
Extended versionhttps://arxiv.org/abs/2511.06899
@@ Introduction
Recent advances in multimodal foundation models have demonstrated increasingly sophisticated capabilities in the combination of visual and textual information  (OpenAI et al. 2024). However, as these models begin to assist in evidentiary reasoning tasks such as criminal case analysis - where establishing reliable connections between surveillance footage (visual modality), forensic reports (textual modality), and other evidence is crucial, and where conclusions must follow rigorous, verifiable reasoning chains - two critical questions emerge: 1. Can current evaluations distinguish between logically valid reasoning and coincidentally correct conclusions? 2. Do existing frameworks capture the non-linear, cross-modal reasoning required to resolve conflicting evidence?
@@
Most existing benchmarks focus solely on task accuracy through multiple choice or short answer formats  (Yu et al. 2024; Yue et al. 2024), completely ignoring the reasoning process. This approach fails to detect when models arrive at correct conclusions through flawed reasoning, a phenomenon we call ”right answers for wrong reasons” , as illustrated on the left side of Figure 1. The few works that examine reasoning processes  (Golovneva et al. 2023; Prasad et al. 2023) typically adopt an oversimplified linear evaluation framework. These approaches are fundamentally mismatched to the complex, non-linear nature of real-world reasoning, where multimodal evidence may appear conflicting yet collectively support valid conclusions. To address these limitations, we introduce a novel evaluation metric: Reasoning Process Tree Score (RPTS), designed to assess multimodal reasoning processes. The core innovation of RPTS lies in its tree-structured representation of reasoning, where leaf nodes correspond to atomic evidence units (visual or textual) and non-leaf nodes capture the hierarchical derivation of intermediate conclusions. This structure inherently accommodates the non-linear interactions characteristic of multimodal reasoning. Furthermore, RPTS incorporates two key hyperparameters whose adjustable values enable precise quantification of both global and local logical consistency, thereby facilitating accurate error localization within the reasoning chain.
@@
Classic multimodal benchmarks typically assess the specific reasoning abilities of the models. For example, OK-VQA(Marino et al. 2019) evaluates a model’s capacity to leverage external knowledge for reasoning, while VCR(Zellers et al. 2019) focuses on human-related common sense reasoning. To evaluate the comprehensive capabilities of a model, researchers have proposed various benchmarks, such as MMBench(Liu et al. 2025), SEED-Bench(Li et al. 2024), MM-Vet(Yu et al. 2024), and MMMU(Yue et al. 2024). These benchmarks scrutinize the reasoning abilities of models from diverse perspectives, often employing multiple choice or simplified formats to facilitate the evaluation process. InfiMM-Eval(Han et al. 2023) incorporates the reasoning process into the evaluation, scoring the entire reasoning process. However, it cannot perform a more detailed analysis of reasoning and its evaluation method cannot exclude cases where incorrect reasoning leads to a correct answer.
@@
Recent studies have introduced various techniques for evaluating reasoning processes. ROSCOE(Golovneva et al. 2023) proposes a set of quality metrics to assess reasoning from four perspectives: semantic alignment, semantic similarity, logical correctness, and semantic coherence. ReCEval (Prasad et al. 2023) evaluates reasoning based on two criteria: whether the reasoning steps are correct and whether new information is derived from the reasoning. REVEAL provides a dataset to validate whether a model can be used to verify the reasoning process.
@@
To construct the reasoning tree, we first parse the model’s reasoning into a structured format: ”[PREMISE] + [PREMISE] + … $\rightarrow$ [CONCLUSION]”, where ’[PREMISE]’ can be derived from visual clues, textual clues, or intermediate conclusions from prior steps. However, existing open-source MLLMs cannot strictly adhere to this output format. To address this, we first employ chain-of-thought (CoT) prompting to guide the model in generating step-by-step reasoning with explicit premises. Subsequently, we use GPT-4 to reformat the reasoning into a structured, easily parsable representation. The parsed reasoning results can also be utilized for accuracy computation.
@@
Now, each reasoning step in our approach strictly adheres to the ”[PREMISE] + [PREMISE] + … $\rightarrow$ [CONCLUSION]” format. Prior studies (Chiang and Lee 2023; Liu et al. 2023; Fu et al. 2024; Bai et al. 2023b; Bitton et al. 2023; Yu et al. 2024; Han et al. 2023) have demonstrated LLMs’ effectiveness in assessing model reasoning. Therefore, we utilize a LLM to score reasoning, but with a unique twist: we only evaluate individual reasoning steps, not the entire process. This method allows for more precise evaluations by preventing the influence of other reasoning elements on the scores. Before we input the reasoning into scorer, we first preprocess the model’s reasoning by eliminating redundant text clues, merging conclusions from images, substituting unnumbered texts and conclusions with all relevant clues and conclusions, and removing reasoning without ’[PERMISE]’. For scoring reasoning according with image, we calculate the semantic similarity of conclusions directly derived from images against the ground truth. For other reasoning, we input the premises and conclusion into LLM to assess their logical coherence. The score given by scorer ranges from 0 to 1, with higher scores indicating stronger logical reasoning. However, as illustrated in Figure 1, there are instances where the model’s selected premises may not directly support the given conclusion, though they may be justified within the broader reasoning context. To address this, if the initial score is below 0.5, we re-evaluate using all text clues and previously derived conclusions as new premises, and then applying a 0.8 penalty for incorrect premises. We select the higher of the two scores as the final assessment.
@@
Figure 5: Three types of relationships between modalities. The filled arrow represents the relation between image and text, and the outlined arrow represents the interference between them.
@@
Figure 4 illustrates the scoring outcomes under three different settings of $\lambda$ and $h_{f}$. If we aim to assign equal importance to each reasoning step, we can set $\lambda$ to 1, as shown in the top-right corner of Figure 4. In this case, RPTS represents the average score of all reasoning steps. Conversely, if we only wish to focus on the first step, i.e., reasoning at height 1, as depicted in the bottom-left corner, we set $\lambda$ to 0 and $h_{f}$ to 1. Under this configuration, only the score of the first step contributes to the RPTS computation. In the main experiment presented in Table 3, the hyperparameters we configured are consistent with those shown in the bottom-right corner of Figure 4. We focused more on the first step in reasoning, so we set $h_{f}=1$. As shown in Table 2, most of the reasoning trees in RPTS-Eval have a height of no more than 7, and we want the weight of each reasoning step to be no less than 0.5 when calculating the RPTS. Based on Formula 1, we can calculate $\lambda>\sqrt[|h_{f}-7|]{0.5}\approx 0.891$. Therefore, we selected $\lambda=0.9$.
@@
To reduce the difficulty of manually designing stories, we use GPT-4 to assist annotators. First, we ask an annotator to design a few reasoning stories and input them into GPT-4 as examples. Following the approach of MM-Vet(Yu et al. 2024), we then require GPT-4 to generate reasoning stories encompassing six types of capabilities, based on the given examples. However, we define two distinct capabilities that differ from MM-Vet: Image Comparison (IC) and Spatial Awareness (SA). The remaining four capabilities—Recognition (Rec), OCR, Commonsense Reasoning (Com), and Math—are identical to those defined in MM-Vet. The specific definitions of IC and SA are as follows:
@@
As mentioned above, the stories generated by GPT-4 lack logical consistency. Therefore, annotators only draw inspiration from these stories rather than using them directly, thereby reducing the difficulty of story design. For example, GPT-4 generates a story set on a rainy day, and our annotators draw inspiration from this, such as the idea of ’rain,’ and design reasoning tasks based on that inspiration. For instance, they might infer whether the box is open or closed based on whether there’s water inside, or they might reason about whether outdoor activities can continue based on certain pre-rain features.
@@
Benchmark Size Images Answer Format Metric Evaluate Reasoning
MMMU(Yue et al. 2024) 11.5K 12.5K Option/Open Answer Accuracy
MM-Vet(Yu et al. 2024) 218 200 Shot answer GPT4-score
InifMM-Eval(Han et al. 2023) 279 342 Reasoning GPT4-score
RPTS-Eval(Ours) 390 374 Reasoning Accuracy+RPTS
@@
To ensure data quality, each piece of data is validated by two validators. We reference InifMM-Eval(Han et al. 2023) and conduct a comprehensive evaluation of the data based on the following criteria:
@@
Experiments
Models and Evaluation Metrics
To validate the challenging nature of RPTS-Eval and the capability of the RPTS evaluation metric analysis model, we conducted experiments in both Chinese and English across various models. The open-source models tested include InstructBLIP(Dai et al. 2024), InternVL2(Chen et al. 2024), ShareGPT4V(Chen et al. 2023), Llava-v1.5(Liu et al. 2024a), Llava-Next(Liu et al. 2024b) and Qwen-VL-Chat(Bai et al. 2023a), detailed in Appendix A; the sole close-source model examined is GPT-4o. We evaluate the reasoning ability of the model by combining accuracy and RPTS, and analyze the problems of the model.
Scorer Selection
Model $\overline{\Delta}$
Qwen2-7B 0.216
Llama-3-8B 0.231
Qwen2-72B 0.143
Llama-3-70B 0.152
GPT-4 0.095
Experiment Settings
Our experiment involves both Chinese and English languages and performs chain-of-thought(COT)(Wei et al. 2022) reasoning on the RPTS-Eval benchmark. All tests were performed in a zero-shot setting using a greedy decoding strategy to assess the models’ inferential abilities. To optimize the COT reasoning outcomes, we designed five Chinese prompts and seven English prompts, selecting the most effective one from each language for our experiments. All tests were carried out on an NVIDIA A100 GPU. When reasoning, we set the temperature of each model to 0 and use greedy decoding.
@@
Table 3 presents model performance on RPTS-Eval. Beyond inference accuracy and mean RPTS scores, we applied an RPTS-based filter (score ¡ 0.5) to exclude cases where correct conclusions arose from flawed reasoning. Results show all models experienced accuracy declines, with GPT-4o least affected—consistent with its stronger logical capacity. In the results of GPT-4, the lower RPTS scores are associated with erroneous reasoning and the model’s failure to capture certain infomation. Conversely, the open-source models demonstrated a lack of logical robustness in their reasoning processes, leading to more pronounced decreases due to often generating irrelevant or illogical outputs. Despite these models’ lower accuracy, their RPTS scores were not significantly impacted. We hypothesize that this is due to two primary reasons: 1. Disconnection between the inference outcomes and the intended targets. While the models initially could reason based on the specified targets, they gradually lost focus on the targets as the number of reasoning steps increased, resulting in conclusions that diverged from the intended data targets. 2. Recurrent generation of identical sentences. Across various sizes, the open-source models consistently produced repetitive reasoning that, while logically sound, failed to reach the desired conclusions. These factors led to reduced accuracy but did not substantially affect the logical integrity of the inferences, as reflected in the relatively high RPTS scores. In addition, Appendix B shows the performance of the model on six capabilities.
@@
V T V T
InternVL2-8B 0.50 0.76 0.54 0.94
Llava-1.5-7B 0.42 0.78 0.40 0.67
Llava-Next-7B 0.36 0.52 0.22 0.58
ShareGPT4V-7B 0.35 0.62 0.30 0.70
Llama-3.2-11B 0.52 0.83 0.4 1.0
InstructBLIP 0.40 0.69 - -
Llava-1.5-13B 0.41 0.66 0.37 0.78
Llava-Next-13B 0.36 0.57 0.29 0.73
Qwen-VL-Chat 0.45 0.74 0.45 0.66
ShareGPT4V-13B 0.18 0.57 0.42 0.75
InternVL2-26B 0.53 0.80 0.57 0.84
Llava-Next-34B 0.54 0.80 0.52 0.72
InternVL2-40B 0.61 0.90 0.60 0.88
InternVL2-76B 0.60 0.92 0.60 0.87
Llama-3.2-90B 0.58 0.79 0.6 0.75
@@ V T V T
Table 5: RPTS score for drawing conclusions from visual clues(V) or textual clues(T).
@@
Table 6: Sensitivity analysis of RPTS: values (top) and percentage of low-score correct answers (bottom) under various $\lambda$ and $h_{f}$.
Step Analysis.
Sensitivity Analysis